# Discrete Diffusion in Large Language and Multimodal Models: A Survey

### v1

@@ II-D Concrete Score
II-D1 Training Loss
Training is typically done by minimizing divergence-based losses that compare the predicted ratio to the true ratio.
Concrete Score Matching (CSM) [29]
The most direct approach is Concrete Score Matching (CSM) [29], which uses a squared-error loss to match the predicted and true ratios:
@@ II-D1 Training Loss
Diffusion-Weighted Denoising Score Entropy [37]
Leveraging Bregman divergence, score entropy is formulated as another score matching loss. Score entropy is non-negative, symmetric, and convex. It is also an extension of the conventional cross-entropy to general positive-valued functions beyond the probability simplex. Score entropy enables the construction of an ELBO for discrete diffusion models, resulting in the Diffusion-Weighted Denoising Score Entropy (DWDSE) loss [37].
@@
$\qquad+K\left(\frac{p_{t|0}(x^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\right)\Bigg{)}\,dt.$ (28)
II-D2 Connection with CE Loss
@@
II-D3 Time Independency
@@
II-E Discrete Flow Matching (DFM)
@@
II-F Block Diffusion Models
@@ III Large Discrete Diffusion Language and Multimodal Models
III-A Discrete Diffusion Models around 1B
Discrete diffusion has emerged as a compelling alternative to continuous diffusion for modeling both discrete data and continuous data. [27] first introduces a diffusion process over binary variables. This idea is generalized to categorical variables by [14], who demonstrates its effectiveness on image generation. Building on these foundations, D3PM [13] proposes a more flexible family of noising schedules that extends discrete diffusion to a broader class of discrete spaces (see Section II-A). DiffusionBERT [32] explores training BERT [73] to reverse a discrete diffusion process with an absorbing state, introducing a token-aware noise schedule for the forward pass and methods to embed time-step information into BERT.
[15] derives an alternative yet equivalent formulation of the sampling from discrete diffusion process (see Section II-B) and introduces a new model family named Reparameterized Discrete Diffusion Models (RDMs). It reformulate the backward process of discrete diffusion in D3PM [13] into a two-stage sampling procedure and yield a greatly simplified training objective and enable more flexible decoding algorithms for text generation. Masked-Diffuse LM (MDLM) [33] propose to leverage the inherit linguistic features of texts to encourage the model to recover the text following an easy-first-generation nature, and directly predict the discrete token with cross-entropy loss to stabilize the intermediate diffusion steps. Diffusion-NAT [34] uses the pretrained BART [74] as the language backbone and unifies the inference process of pretrained language models and the denoising process of discrete diffusion models in a non-autoregressive manner, thus the BART model plays the role of the parameterized denoiser in discrete diffusion models. Furthermore, TESS [35] leverages a new form of self-conditioning and applies diffusion on the logit simplex instead of the learned embedding space. Plaid [36] takes the first step towards closing the likelihood gap between autoregressive and diffusion-based language models through algorithmic improvements, scaling laws, and increased compute. SEDD [37] generalize the idea of score matching into the discrete spaces by proposing a novel loss named score entropy, which can be integrated seamlessly to build discrete diffusion models and can significantly boost model performance.
@@ III-A Discrete Diffusion Models around 1B
III-B Large Diffusion Language Models
III-B1 LLaDA
@@ III-B6 TESS 2
III-C Large Diffusion Multimodal Models
III-C1 Dimple
@@
During inference, Dimple introduces a confident decoding strategy for efficiency: the model dynamically chooses how many tokens to fill in at each step based on model confidence (see Sec. V-A). Empirically, this reduces the number of iterations to about response $\frac{\text{response length}}{3}$. Dimple also re-implements an autoregressive prefilling trick: by filling in some tokens from the existing context, it speeds up inference by about 1.5$\times$–7$\times$ with minimal impact on quality. Under the same training budget and dataset as LLaVA-NEXT [89], Dimple-7B achieves higher aggregate scores on multimodal benchmarks than LLaVA-NEXT-7B. This result shows that with a proper hybrid training recipe, a discrete dMLLM can match or exceed strong autoregressive baselines on vision-language tasks.
III-C2 LaViDa
@@ III-C3 LLaDA-V
III-D Large Unified Model
III-D1 MMaDA
@@ IV-A1 Low Corpus Utilization
IV-A2 Random Sampling of Time Index
@@ IV-B1 BERT Initialization
IV-B2 Autoregressive Model Initialization
@@
IV-B3 Autoregressive-then-Diffusion Training
@@
IV-C Complementary Masking Technique
@@
IV-D Masking Scheduling Technique
@@
IV-D1 Uniform Masking Scheduling
@@
where $\alpha_{t}^{\prime}=\frac{d\alpha_{t}}{dt}$ is the derivative of $\alpha_{t}$ with respect to time. Several scheduling strategies have been proposed and empirically studied:
@@
    $w_{t}$ $=\left(\frac{\exp\left(-\bar{\beta}_{\min}^{1-t}\bar{\beta}_{\max}^{t}\right)}{1-\exp\left(-\bar{\beta}_{\min}^{1-t}\bar{\beta}_{\max}^{t}\right)}\right)\bar{\beta}_{\min}^{1-t}\bar{\beta}_{\max}^{t}\log\left(\frac{\sigma_{\min}}{\sigma_{\max}}\right).$ (50)
@@
IV-D2 Token-wise Masking Scheduling
@@ V-A Unmasking Techniques
V-A1 Discrete-Time Unmasking
@@
    indicating the model’s certainty about the most likely token.
@@
V-A2 Continuous-Time Unmasking (Flow Matching)
@@ V-B Remasking Techniques
Remasking in General Discrete Diffusion Models.
@@
Remasking under Discrete Flow Matching.
@@
V-C Prefilling and Caching Techniques
@@ V-C1 Prefilling
V-C2 KV-Cache
Due to the use of bidirectional attention, the cached KV pairs, attention outputs, and other values in dLLMs and dMLLMs are not static. The associated caching algorithms typically consist of three components: caching, reuse, and update. The papers dKV-Cache [25] and dLLM-Cache [26] were among the first to introduce caching techniques into dLLMs.

### v2

@@ II.E Concrete Score
II.E.1 Training Loss
[…]
@@ II.E.1 Training Loss
[…]
Diffusion-Weighted Denoising Score Entropy [40] Leveraging Bregman divergence, score entropy is formulated as another score matching loss. Score entropy is non-negative, symmetric, and convex. It is also an extension of the conventional cross-entropy to general positive-valued functions beyond the probability simplex. Score entropy enables the construction of an ELBO for discrete diffusion models, resulting in the Diffusion-Weighted Denoising Score Entropy (DWDSE) loss [40]
@@
$\qquad+K\left(\frac{p_{t|0}(x^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\right)\Bigg{)}\,dt,$ (38)
where $K$ is a normalizing constant function.
[…]
Target Concrete Score Matching [33] Target Concrete Score Matching (TCSM) introduces two score matching losses: the score-based and distribution-based losses. The score-based objective operates directly on the concrete score vectors:
$\mathcal{L}_{\mathrm{score}}(\theta)=\mathbb{E}_{\omega(t)p(x_{t})h(x_{1}|x_{t})}\sum_{i=1}^{L}\ell^{i}_{\text{score}},$ (39)
$\ell^{i}_{\text{score}}=\mathcal{D}\left(\left[\frac{p_{1|t}([y^{i},x_{1}^{\backslash i}]|x_{t})}{p_{1|t}(x_{1}|x_{t})}\right]_{y^{i}=1}^{V},\left[\frac{p^{\theta}_{1|t}([y^{i},x_{1}^{\backslash i}]|x_{t})}{p_{\theta}^{1|t}(x_{1}|x_{t})}\right]_{y^{i}=1}^{V}\right),$ (40)
where $V$ is the vocabulary size, $\mathcal{D}(\cdot,\cdot)$ denotes a divergence measure, $\omega(t)$ is the distribution used to sample time index $t$, and $h(x_{1}|x_{t})$ is a proposal distribution such as the ground-truth conditional distribution $p_{1|t}(x_{1}|x_{t})$. In the equation, $[y^{i},x^{\backslash i}]:=[x^{1},\ldots,x^{i-1},y^{i},x^{i+1},\ldots,x^{L}]$ is used to define a new sequence of the same length as $x_{1}$, where the token at position $i$ is replaced by $y^{i}$ and all other tokens remain identical to those in $x_{1}$. $x^{\backslash i}_{1}$ is used to indicate all tokens in $x_{1}$ except for the token at position $i$.
The distribution-based objective aligns the model and true conditional distributions:
$\mathcal{L}_{\mathrm{distrib}}(\theta)=\mathbb{E}_{\omega(t)p(x_{t})}\sum_{i=1}^{L}\mathbb{E}_{h(x_{1}^{\backslash i}|x_{t})}\ell^{i}_{\text{distrib}},$ (41)
$\ell^{i}_{\text{distrib}}=\mathbb{D}\left(p_{1|t}(x_{1}^{i}|x_{1}^{\backslash i},x_{t})\,\|\,p^{\theta}_{1|t}(x_{1}^{i}|x_{1}^{\backslash i},x_{t})\right),$ (42)
where $\mathbb{D}(\cdot,\cdot)$ is a statistical divergence measures the differences between probability distributions. Shown in Proposition 3 of [33], with $h(x_{1}|x_{t})=p^{1|t}(x_{1}|x_{t})$, the two objectives are equivalent:
$\mathcal{L}_{\mathrm{score}}(\theta;\mathcal{D}=\mathcal{D}_{\mathrm{GKL}})\equiv\mathcal{L}_{\mathrm{distrib}}(\theta;\mathbb{D}=V\mathbb{D}_{\mathrm{KL}}+\mathbb{D}_{\mathrm{IS}}),$ (43)
where $\mathcal{D}_{\mathrm{GKL}}$, $\mathbb{D}_{\mathrm{KL}}$ and $\mathbb{D}_{\mathrm{IS}}$ refer to the generalized Kullback–Leibler divergence, the Kullback–Leibler divergence and Itakura–Saito divergence, respectively.
By selecting different discrepancy measures and proposal distributions, there are different instantiations of the TCSM loss. For instance, choosing the generalized KL divergence as the discrepancy $\mathcal{D}$ and the true conditional distribution $p_{1|t}(x_{1}|x_{t})$ as the proposal $h(x_{1}|x_{t})$, the score-based TCSM becomes
$\ell_{\mathrm{score}}^{i}$ $=\left(-\log p^{\theta}_{1|t}(x_{1}^{i}|x_{t})+\frac{1}{Vp^{\theta}_{1|t}(x_{1}^{i}|x_{t})}\right)$
$\qquad+\frac{1}{V}\sum_{y}\log p^{\theta}_{1|t}(y|x_{t});$ (44)
choosing the KL divergence as the discrepancy $\mathbb{D}$ and the true conditional distribution $p_{1|t}(x_{1}|x_{t})$ as the proposal $h(x_{1}|x_{t})$, the distribution-based TCSM becomes
$\ell_{\mathrm{distrib}}^{i}=-\mathbb{E}_{p_{1|t}}\log p^{\theta}_{1|t}(x_{1}^{i}|x_{t})+C,$ (45)
where $C$ is a constant. The distribution-based objective reduces to a cross-entropy loss, maximizing the pseudo-likelihood of $p^{\theta}_{1|t}$ under the true denoising distribution.
II.E.2 Connection with CE Loss
@@
[28] also connects the concrete score matching with the CE loss. Consider two discrete sequences $x$ and $y$ that only differ at position $i$. By Bayesian Theorem, the probability ratio $\frac{p_{t}(y)}{p_{t}(x)}$ can be rewritten as
$\frac{p_{t}(y)}{p_{t}(x)}=\sum_{h\in\mathcal{V}}\frac{p_{t|0}(y^{i}\mid h)}{p_{t|0}(x^{i}\mid h)}\cdot p^{i}_{0|t}(h\mid x_{t}),$ (47)
where $p_{t|0}(\cdot\mid h)$ is the known conditional transition probability in the forward process, and $p^{i}_{0|t}(h\mid x_{t})$ is the posterior of the clean token $h$ at position $i$ given the noised sequence $x_{t}$. Thus, the concrete score can be parameterized as
$s^{i}_{\theta}(x_{t},t)=\sum_{h\in\mathcal{V}}\frac{p_{t|0}(y^{i}\mid h)}{p_{t|0}(x^{i}\mid h)}\cdot f^{i}_{\theta}(x_{t},t)[h],$ (48)
where $f^{i}_{\theta}(x_{t},t)[h]$ models the likelihood that token $h$ was the original clean token at position $i$. This allows training to proceed via a simple cross-entropy loss over the posterior $p_{0|t}$, rather than requiring explicit score estimation:
$\mathcal{L}_{\text{CEDD}}=\mathbb{E}_{t}\mathbb{E}_{x_{0},x_{t}}\sum_{i=1}^{L}w(t)\cdot\log f^{i}_{\theta}(x_{t},t)[x^{i}_{0}],$ (49)
where $w(t)$ is a timestep-dependent weighting function.
II.E.3 Time Independency
@@
II.F Discrete Flow Matching (DFM)
@@
II.G Block Diffusion Models
@@ III Large Discrete Diffusion Language and Multimodal Models
III.A Discrete Diffusion Models around 1B
Discrete diffusion has emerged as a compelling alternative to continuous diffusion for modeling both discrete data and continuous data. [27] first introduces a diffusion process over binary variables. This idea is generalized to categorical variables by [14], who demonstrates its effectiveness on image generation. Building on these foundations, D3PM [13] proposes a more flexible family of noising schedules that extends discrete diffusion to a broader class of discrete spaces (see Section II.A). DiffusionBERT [35] explores training BERT [81] to reverse a discrete diffusion process with an absorbing state, introducing a token-aware noise schedule for the forward pass and methods to embed time-step information into BERT.
[15] derives an alternative yet equivalent formulation of the sampling from discrete diffusion process (see Section II.B) and introduces a new model family named Reparameterized Discrete Diffusion Models (RDMs). It reformulate the backward process of discrete diffusion in D3PM [13] into a two-stage sampling procedure and yield a greatly simplified training objective and enable more flexible decoding algorithms for text generation. Masked-Diffuse LM (MDLM) [36] propose to leverage the inherit linguistic features of texts to encourage the model to recover the text following an easy-first-generation nature, and directly predict the discrete token with cross-entropy loss to stabilize the intermediate diffusion steps. Diffusion-NAT [37] uses the pretrained BART [82] as the language backbone and unifies the inference process of pretrained language models and the denoising process of discrete diffusion models in a non-autoregressive manner, thus the BART model plays the role of the parameterized denoiser in discrete diffusion models. Furthermore, TESS [38] leverages a new form of self-conditioning and applies diffusion on the logit simplex instead of the learned embedding space. Plaid [39] takes the first step towards closing the likelihood gap between autoregressive and diffusion-based language models through algorithmic improvements, scaling laws, and increased compute. SEDD [40] generalize the idea of score matching into the discrete spaces by proposing a novel loss named score entropy, which can be integrated seamlessly to build discrete diffusion models and can significantly boost model performance.
@@ III.A Discrete Diffusion Models around 1B
III.B Large Diffusion Language Models
III.B.1 LLaDA
@@ III.B.6 TESS 2
III.C Large Diffusion Multimodal Models
III.C.1 Dimple
@@
During inference, Dimple introduces a confident decoding strategy for efficiency: the model dynamically chooses how many tokens to fill in at each step based on model confidence (see Sec. V.A). Empirically, this reduces the number of iterations to about response $\frac{\text{response length}}{3}$. Dimple also re-implements an autoregressive prefilling trick: by filling in some tokens from the existing context, it speeds up inference by about 1.5$\times$–7$\times$ with minimal impact on quality. Under the same training budget and dataset as LLaVA-NEXT [97], Dimple-7B achieves higher aggregate scores on multimodal benchmarks than LLaVA-NEXT-7B. This result shows that with a proper hybrid training recipe, a discrete dMLLM can match or exceed strong autoregressive baselines on vision-language tasks.
III.C.2 LaViDa
@@ III.C.3 LLaDA-V
III.D Large Unified Model
III.D.1 MMaDA
@@ IV.A.1 Low Corpus Utilization
IV.A.2 Random Sampling of Time Index
@@ IV.B.1 BERT Initialization
IV.B.2 Autoregressive Model Initialization
@@
IV.B.3 Autoregressive-then-Diffusion Training
@@
IV.C Complementary Masking Technique
@@
IV.D Masking Scheduling Technique
@@
IV.D.1 Uniform Masking Scheduling
@@
where $\alpha_{t}^{\prime}=\frac{d\alpha_{t}}{dt}$ is the derivative of $\alpha_{t}$ with respect to time. Several scheduling strategies have been proposed and empirically studied. The corresponding schedules are plotted in Fig. 4.
@@
    $w_{t}$ $=\left(\frac{\exp\left(-\bar{\beta}_{\min}^{1-t}\bar{\beta}_{\max}^{t}\right)}{1-\exp\left(-\bar{\beta}_{\min}^{1-t}\bar{\beta}_{\max}^{t}\right)}\right)\bar{\beta}_{\min}^{1-t}\bar{\beta}_{\max}^{t}\log\left(\frac{\bar{\beta}_{\min}}{\bar{\beta}_{\max}}\right).$ (70)
@@
  • Polynomial Schedule [17]:
    $\alpha_{t}$ $=1-t^{r},$ (73)
    $w_{t}$ $=-\frac{r}{t}.$ (74)
IV.D.2 Token-wise Masking Scheduling
@@
(a) $\alpha_{t}$ Schedules
(b) $w_{t}$ Schedules
Figure 4: Different schedules for $\alpha_{t}$ and $w_{t}$. To unify notation, we also transformed the spindle schedule from the discrete-time format used in the original paper to a continuous-time format. The revised formulation is as follows: $\alpha_{t}=1-t-\tau\sin(\pi t)$ and $w_{t}=-\frac{1+\tau\pi\cos(\pi t)}{t+\tau\sin(\pi t)}$, where $\tau$ corresponds to the original $\lambda\tilde{H}$.
IV.E Addressing Training-Testing Discrepancy
[48] mentioned a discrepancy between training and inference of dLLM. During training, the model receives ground-truth noisy tokens as input, while at inference time, the inputs of the model are the previously predicted tokens. To address this, [48] propose the two-step loss.
Let $x_{0}$ denote the ground truth sequence. During training, a time step $t$ is randomly selected, and a noised version $x_{t}$ is generated from $x_{0}$ via a diffusion process. The model then predicts the original sequence $\hat{x}_{0}=f_{\theta}(x_{t})$. Subsequently, a second input $\hat{x}_{t-1}$ is generated by applying the forward diffusion transition matrixto the predicted sequence $\hat{x}_{0}$ and the model again attempts to recover the ground truth $\hat{\hat{x}}_{0}=f_{\theta}(\hat{x}_{t-1})$. The two-step loss is calculated between the ground truth $x_{0}$ and the twice-denoised output $\hat{\hat{x}}_{0}$.
To ease training in early stages, the model does not always use this two-step strategy. Instead, a mixed strategy is adopted. With probability $1-p_{k}$, the two-step strategy is used and the loss is evaluated between $\hat{\hat{x}}_{0}$ and $x_{0}$. With probability $p_{k}$, the conventional one-step strategy is used and the loss is evaluated between $\hat{x}_{0}$ and $x_{0}$. $p_{k}$ is set to be linearly increasing along the training step $k$.
IV.F Reweighting Technique
IV.F.1 Multi-Granularity Diffusion Modeling
Multi-Granularity Diffusion Modeling (MGDM) [49] introduces an additional token-level reweighting factor $v(x_{t,n})$, yielding a refined training loss:
$\mathcal{L}_{\text{MGDM}}=\sum_{n=1}^{N}\sum_{t=1}^{T}w(t)\cdot v(x_{t,n})\cdot\ell(x_{0},x_{t},n;\theta),$ (78)
where $\ell(x_{0},x_{t},n;\theta)$ is the CE loss on the n-th token, and the adaptive token-level weight is defined as:
$v(x_{t,n})=\alpha(1-\exp(-\ell(x_{0},x_{t},n;\theta)))^{\beta},$ (79)
with hyperparameters $\alpha>0$, $\beta>0$. This reweighting assigns larger weights to harder tokens (i.e., those with higher loss), thereby prioritizing difficult subgoals during training and accelerating convergence.
IV.G Distillation through Dimensional Correlations
To enable efficient few-step or even one-step generation while preserving performance, Di4C [50] explicitly models inter-dimensional correlations and introduces two principled loss functions: Distillation Loss and Consistency Loss. These losses are grounded in distributional matching and multi-path coherence.
IV.G.1 Distillation Loss
The distillation loss transfers the probabilistic knowledge from a teacher model, which performs multi-step denoising, to a student model that aims to generate outputs in fewer steps. This is achieved by aligning the student’s predicted posterior with that of the teacher at an intermediate noise level $\delta$. Formally, the loss is defined as:
$\mathcal{L}_{\text{distil}}(\theta;\psi,r_{\delta},\delta)$ $=\mathbb{E}_{x_{\delta}\sim r_{\delta}}\Big{[}\mathrm{D_{KL}}\big{(}p^{\psi}_{0|\delta}(\cdot\mid x_{\delta})$
$\qquad\|\,p^{\theta}_{0|\delta}(\cdot\mid x_{\delta})\big{)}\Big{]}$ (80)
where $p^{\psi}_{0|\delta}(\cdot|x_{\delta})$ is the posterior distribution over clean data $x_{0}$ given intermediate noisy input $x_{\delta}$ under the teacher model; $p^{\theta}_{0|\delta}(\cdot|x_{\delta})$ is the student model’s predicted posterior; $r_{\delta}$ is a reference distribution over noisy states (typically chosen to match the forward diffusion at timestep $\delta$). This loss encourages to transfer the full-step generative knowledge from teacher to student by matching posteriors.
IV.G.2 Consistency Loss
The consistency loss ensures that the student model behaves consistently when predicting from different intermediate noise levels. Specifically, if $x_{t}$ is a noisy sample at step $t$, there should be agreement between: 1). First denoising $x_{t}\to x_{u}$ via the teacher, then $x_{u}\to x_{s}$ via the student; 2). Directly predicting $x_{t}\to x_{s}$ via the student. This intuition is captured by the following KL divergence:
$\mathcal{L}_{\text{consis}}(\theta;\psi,r_{t},s,u,t)={}$ $\mathbb{E}_{x_{t}\sim r_{t}}\Big{[}\mathrm{D_{KL}}\bigl{(}(p^{\theta}_{s|u}\circ p^{\psi}_{u|t})(\cdot\mid x_{t})$
$\|\,p^{\theta}_{s|t}(\cdot\mid x_{t})\bigr{)}\Big{]}$ (81)
where $p^{\psi}_{u|t}(x_{u}|x_{t})$ is the teacher’s distribution from timestep $t$ to $u$; $p^{\theta}_{s|u}(x_{s}|x_{u})$ is the student’s distribution from $u$ to $s$; $p^{\theta}_{s|t}(x_{s}|x_{t})$ is the direct prediction by the student; $(p^{\theta}_{s|u}\circ p^{\psi}_{u|t})(\cdot|x_{t})$ denotes the composite distribution over $x_{s}$ via intermediate $x_{u}$. This loss enforces functional coherence across generation paths, capturing multi-dimensional correlations without assuming independence.
@@ V.A Unmasking Techniques
V.A.1 Discrete-Time Unmasking
[…]
@@ V.A.1 Discrete-Time Unmasking
    indicating the model’s certainty about the most likely token. [51] provides a theoretical analysis of the equivalence between the parallel decoding using confidence and sequential decoding using confidence.
@@
V.A.2 Continuous-Time Unmasking (Flow Matching)
@@ V.B Remasking Techniques
V.B.1 Remasking in General Masked Diffusion Models.
@@
V.B.2 Remasking under Discrete Flow Matching.
@@
V.C Prefilling and Caching Techniques
@@ V.C.1 Prefilling
V.C.2 KV-Cache
Due to the use of bidirectional attention, the cached KV pairs, attention outputs, and other values in dLLMs and dMLLMs are not static. The associated caching algorithms typically consist of three components: caching, reuse, and update.